OpenStudy (anonymous):

A rectangular garden has a perimeter of 172 feet. The length of the garden is 24 feet more than the width. What are the dimensions of the garden?

OpenStudy (anonymous):

alright, first I think you would do 24+24 which is 48, then subtract that from 172, and you get 124, divide that by 2 and you should get your answer I thnk

OpenStudy (anonymous):

let l = length and let w=width. l=w+24 (saying that it order for width to be equal to length, you would add 24 more to the width. The width is shorter). P=2(w+l) 172=2(w+l) 172=2(w+w+24) 172=2(2w+24) 86=(2w+24) 86=2w+24 62=2w w=31 so, l=w+24 l=31+24=55 SO 55 and 31, Checking.... 2(55+31)=172 2(66)=172 172=172